OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (anonymous):

@Directrix please help thank you

OpenStudy (anonymous):

I think the answer is C x - y = - 5

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

the points (-2, 3) and (2, 7)?

OpenStudy (anonymous):

you are right you got it

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

yep I am sure

OpenStudy (anonymous):

@Directrix what would you say

Directrix (directrix):

I got x - y + 5 = 0 which is another way to write x - y = -5.

OpenStudy (anonymous):

ok thank you

Directrix (directrix):

You can use this online calculator of a line. Enter the two points and crank out the equation. http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

OpenStudy (anonymous):

thats pretty cool how'd you find out about this

OpenStudy (anonymous):

I smart

Directrix (directrix):

Googled it up.

OpenStudy (anonymous):

yeah he is right google it